OpenStudy (twizttiez):

Solution or no Solution? 4(1−x)=5x+8

OpenStudy (alexandervonhumboldt2):

4-4x=5x+8 -9x=4 so there is a solution

OpenStudy (alexandervonhumboldt2):

x=-4/9

OpenStudy (twizttiez):

I have one more can you help?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (twizttiez):

Thanks x+5+2x=5+3x

OpenStudy (alexandervonhumboldt2):

3x-3x=5-5 0=0 there is an infinite number of solutions

OpenStudy (twizttiez):

Thanks so much if i need help can i just tag you?

OpenStudy (alexandervonhumboldt2):

yes tyag me

OpenStudy (twizttiez):

Ok cool

OpenStudy (twizttiez):

2. A skydiver jumps from a plane from an altitude of 6000 ft. An observer on the ground measures the skydiver’s altitude at time t = 0 to be 4500 ft. The linear graph shows the skydiver’s altitude at various times during her descent.

OpenStudy (twizttiez):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

what to find?

OpenStudy (twizttiez):

Do you understand it i dont

OpenStudy (twizttiez):

(a) Identify the y-intercept. How does the y-intercept relate to the skydiver’s position relative to the ground? (b) Determine the slope of the line. According to the slope, is the skydiver ascending or descending, and what is the rate of ascent or descent? (c) Identify the x-intercept. What does the x-intercept mean in this context?

OpenStudy (twizttiez):

When you answer each one put the letter next to which question you answered

OpenStudy (twizttiez):

I hate to sound bossy btw :|

OpenStudy (twizttiez):

I just need this done asap

OpenStudy (twizttiez):

@Reaper534

OpenStudy (anonymous):

@SolomonZelman Could you help Im not familiar with this and trying to finish this book report thingy xD